OpenStudy (anonymous):

Use factoring to find the zeros of each equation y=x2+2x-24 Answers: A.4,6 B.4,-6 C.-4,-6 D.-4,6

OpenStudy (anonymous):

B

OpenStudy (usukidoll):

if we can not give out direct answers that would be great! ... At least show it's really B for the equation \[y=x^2+2x-24\] notice that the middle term is 2 and the last term is -24 so we need to find what two numbers we need to produce a positive 2

OpenStudy (usukidoll):

all possible combinations of 24 are 1 24 24 1 12 2 2 12 8 3 3 8 4 6 6 4 thank goodness for multiple choice, so we can use the 4 6 6 4 combinations

OpenStudy (usukidoll):

since or equation patter is a + and - we need to figure out which number has to be negative and which number has to be positive in order to achieve 2x and - 24

OpenStudy (usukidoll):

*our equation pattern... sorry typing fast \[y=x^2+2x-24\]

OpenStudy (usukidoll):

I think that -4 and 6 will be perfect.. because of 2 reasons.. we know that (-4)(6) = -24 and -4+6 = 2 so now by factoring we have \[y=(x-4)(x+6)\] we can check using the FOIL method and it does come back to the original equation that we had earlier.

OpenStudy (usukidoll):

to find our 0's we can just let y be 0 and solve for x. I think you could handle it from here 0=(x-4)(x+6) It's like trying to ask yourself what x value do I need for this to hold true.

OpenStudy (usukidoll):

@Caylahill do you want me to go further or can you handle 0=(x-4)(x+6) and find all x's ?

OpenStudy (usukidoll):

you can just use inspection and see that there are 2 x values that will make the equation 0